OpenStudy (anonymous):

e^(1+3t)^2 Chain rule thanks

OpenStudy (anonymous):

the question is [e^(1+3t)]^2 ? if so: 2 * e^(1+3t) * (3)

OpenStudy (anonymous):

which is 6e^(1+3t)

OpenStudy (anonymous):

yea thats what I thought answer in the back is 6(1+3t)e^1+3t)^2

OpenStudy (anonymous):

mm maybe they meant : e^[(1+3t)^2] differentiating this : [(1+3t)^2]' * e^[(1+3t)^2] = 2(1+3t) * 3 * e^[(1+3t)^2] = 6 (1+3t) * e^[(1+3t)^2]

OpenStudy (anonymous):

which is what they got .. so it depends on the problem :)

OpenStudy (anonymous):

ok thats it thanks did I type it wrong?

OpenStudy (anonymous):

this : [e^(1+3t)]^2 is different from : e^[(1+3t)^2]

OpenStudy (anonymous):

maybe i just misunderstood ..

OpenStudy (anonymous):

what are the rules in order I dont get why Im doing what Im doing

OpenStudy (anonymous):

or I guess why did the 1+3t come down in front, did I use the power rule first?